OpenStudy (anonymous):

Find all the rational roots of 2x^3-x^2+8x-4=0

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it really says "all rational roots" right not "all POSSIBLE rational root"?

OpenStudy (anonymous):

yeah it sais "all" :)

OpenStudy (misty1212):

because the only rational root is \(\frac{1}{2}\) as it factors as \[(2x-1)(x^2+4)\]

OpenStudy (misty1212):

if they left out the word "possible" then they want the actual one, namely \(\frac{1}{2}\)

OpenStudy (anonymous):

wait so is it just 1/2? im very bad at algebra haha

OpenStudy (misty1212):

yeah it is just \(\frac{1}{2}\)

OpenStudy (anonymous):

oh okay then thank you!

OpenStudy (misty1212):

unless they said find all "possible' ones if they did not use the word "possible' then only one half

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]